# Leveraging Reinforcement Learning, Genetic Algorithms and Transformers for background determination in particle physics

### v1

@@ Appendix B GA space size
This appendix presents the different types of decay structures of the backgrounds for an example signal within the toy scenario introduced in Section 4, namely , which corresponds to the decay of a positively charged mother particle into two positively charged, one negatively charged, and two neutral final-state particles. Table LABEL:space_size_chart organises these structures according to the number of intermediate resonances (IRs), the number of non-reconstructed particles, the charge of the mother particle, and the charges of the IRs present in the background. The GA space size of each category is also included. While we have attempted to enumerate all possible cases, it cannot be excluded that additional subcases exist. Nevertheless, the list of decay structures provided here is intended to be representative rather than exhaustive. The computation of the total GA space size for this example signal yields 333,305,600101010The computed size represents the search space that the GA can explore, not the number of physically allowed background processes..

### v2

@@ Appendix B GA space size
This appendix presents the different types of decay structures of the backgrounds for an example signal within the toy scenario introduced in Section 4, namely , which corresponds to the decay of a positively charged mother particle into two positively charged, one negatively charged, and two neutral final-state particles. Table LABEL:space_size_chart organises these structures according to the number of intermediate resonances (IRs), the number of non-reconstructed particles, the charge of the mother particle, and the charges of the IRs present in the background. The GA space size of each category is also included. While we have attempted to enumerate all possible cases, it cannot be excluded that additional subcases exist. Nevertheless, the list of decay structures provided here is intended to be representative rather than exhaustive. The computation of the total GA space size for this example signal yields 333,305,600111111The computed size represents the search space that the GA can explore, not the number of physically allowed background processes..